OpenStudy (anonymous):

which of the following is not a solution of the equation 3x-4y=8? A.(0,-2) B.(4,5) C.(3,.25) D.(-4,-1)

OpenStudy (anonymous):

plug each ordered pair into the equation 3( )-4( )=8. For example with the first ordered pair, (0,-2), x=0 and y=-2 so the equation becomes 3(0)-4(-2)=8 0 + 8 = 8 which is a true statement so (0,-2) is a solution of the equation 3x-4y=8 if you get something weird like 7 = 8 when you plug an ordered pair in, then that ordered pair is not a solution.

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

Let's see, for c we have (3, .25) so x = 3 and y = .25 Plugging them in we get 3(3)-4(.25)=8 which is 9 - 1 = 8 right? That's a true equation so (3, .25) is a solution. It isn't c.